Calculate the consumer and producer surpluses before and after a tax is imposed on the production of the following good. Also calculate the tax revenue, DWL, and percent of incidence on each side of the market. 4) Price S+$5 Tax CS Pre-Tax PS Pre-Tax CS Post-Tax PS Post-Tax Tax Revenue DWL 20 18 15 Incidence: % on Consumers % on Producers 0 50 60 Quantity (#/Time)
Answer #1

CS pre tax = (1/2)*(30 - 18)*60 = 360

PS pre tax = (1/2)*18*60 = 540

CS post tax = (1/2)*(30 - 20)*50 = 250

PS post tax =(1/2)*(15 - 0)*50 = 375

Tax revenue = 5*50 = 250

DWL = (1/2)*(20 - 15)*(60 - 50 ) = 25

Incidence :

% on consumers = 20 - 18 = 2

so, 2/5 = 0.4 = 40%

% on producers = 3/5 =0.6 = 60%